# Inertial Coordination Games

### v2

@@ Appendix IV Finite Players
$u(a_{it},\lambda_{N,-i,t-1},\theta):=\begin{cases}\theta+\lambda_{N,-i,t-1}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
@@
is the empirical proportion of non-$i$ agents who took the risky action at time $t-1$ for $t>1$, and $\lambda_{N,-i,t-1}:=\lambda_{0}$ is common knowledge. Let
@@
Inductive step: Fix $t>1$ and assume that at time $t-1$, each agent $i$ takes the risky action iff $\mu_{i,t-1}\geq\mu_{t-1}^{*}$. At time $t$, each agent $i$ takes the risky action iff
@@
[…]
@@ Appendix V Alternative Interpretations of Model
Overlapping generations model. There is an initial unit mass of agents with proportion $\lambda_{0}\in(0,1)$ who take the risky action, all of whom share a common improper uniform prior about $\theta$. At each time $t\geq 1$, a unit mass of children is born (generation $t$), each of whom inherits her prior about $\theta$ from her parent’s (generation $t-1$) belief. Each child $i$ receives a signal $x_{it}\sim N(\theta,\sigma_{t}^{2})$, updates her belief about $\theta$, chooses an action, and gets uniformly matched to play the static game below with a member of generation $t-1$. Finally, child $i$ becomes parent $i$ and continues to take her time-$t$ action whenever matched with an agent from generation $t+1$. We interpret this as due to switching costs—since agents’ actions when playing against the future generation are not payoff-relevant, switching costs ensure this.
@@
with $\mu_{1}^{*}=c-\lambda_{0}$. As we have shown, this equation has a unique fixed point at $c-(1/2)$ and hence we have global convergence $\mu_{t}^{*}\to c-(1/2)$ for any $\lambda_{0}\in(0,1)$. Furthermore, the entire path of $(\mu_{t}^{*})_{t}$ corresponds to the iterated best-response dynamics with symmetric switching strategies in Carlsson and Van Damme (1993)’s two-player setting with the informational environment from above, in which $\mu_{\infty}^{*}$ is the unique rationalizable switching strategy. To see this, let $\mathbb{P}_{x}(\theta,x^{\prime})$ denote the joint posterior over the state $\theta$ and my opponent’s signal $x^{\prime}$, conditional on my signal $x$.
@@
[…]
@@
An analogous argument allows us to extend the interpretation of our model dynamics as two-player iterated joint learning and best-response dynamics by considering our model for arbitrary $\Sigma$. In this case, note that if my opponent receives signal distributed $N(\theta,\eta_{t-1}^{2})$ (corresponding to the time-$(t-1)$ distribution of posterior means, conditional on the state) and my belief is about the state is $N(\mu_{it},\eta_{t}^{2})$, then my (state-unconditional) belief about my opponent’s signal is
@@ Appendix VI Payoffs generalizations
$u(a_{it},\lambda_{t-1},\theta):=\begin{cases}a\theta+b\lambda_{t-1}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$

### v3

@@ Appendix IV Finite Players
[…]
@@
is the empirical proportion of non-$i$ agents who took the risky action at time t1t-1 for $t>1$, and $\lambda_{N,-i,t-1}:=\lambda_{0}$ is common knowledge. Let
@@
Inductive step: Fix $t>1$ and assume that at time t1t-1, each agent $i$ takes the risky action iff $\mu_{i,t-1}\geq\mu_{t-1}^{*}$. At time $t$, each agent $i$ takes the risky action iff
@@
$|\lambda_{N,\infty}-\lambda_{\infty}(\theta)|=0\quad\mathbb{P}_{\theta}-\text{a.s.}$
@@ Appendix V Alternative Interpretations of Model
Overlapping generations model. There is an initial unit mass of agents with proportion $\lambda_{0}\in(0,1)$ who take the risky action, all of whom share a common improper uniform prior about $\theta$. At each time $t\geq 1$, a unit mass of children is born (generation $t$), each of whom inherits her prior about $\theta$ from her parent’s (generation t1t-1) belief. Each child $i$ receives a signal $x_{it}\sim N(\theta,\sigma_{t}^{2})$, updates her belief about $\theta$, chooses an action, and gets uniformly matched to play the static game below with a member of generation t1t-1. Finally, child $i$ becomes parent $i$ and continues to take her time-$t$ action whenever matched with an agent from generation $t+1$. We interpret this as due to switching costs—since agents’ actions when playing against the future generation are not payoff-relevant, switching costs ensure this.
@@
with $\mu_{1}^{*}=c-\lambda_{0}$. As we have shown, this equation has a unique fixed point at c(1/2)c-(1/2) and hence we have global convergence $\mu_{t}^{*}\to c-(1/2)$ for any $\lambda_{0}\in(0,1)$. Furthermore, the entire path of $(\mu_{t}^{*})_{t}$ corresponds to the iterated best-response dynamics with symmetric switching strategies in Carlsson and Van Damme (1993)’s two-player setting with the informational environment from above, in which $\mu_{\infty}^{*}$ is the unique rationalizable switching strategy. To see this, let $\mathbb{P}_{x}(\theta,x^{\prime})$ denote the joint posterior over the state $\theta$ and my opponent’s signal $x^{\prime}$, conditional on my signal $x$.
@@
$=\mathbb{E}_{x}[\theta]+\mathbb{P}_{x}(x^{\prime}\geq\mu^{\prime})=x+\int_{\theta}\mathbb{P}(x^{\prime}\geq\mu^{\prime}|\theta,x)d\mathbb{P}_{x}(\theta)=x+\mathbb{E}_{x}[\lambda(\theta)]$
@@
An analogous argument allows us to extend the interpretation of our model dynamics as two-player iterated joint learning and best-response dynamics by considering our model for arbitrary $\Sigma$. In this case, note that if my opponent receives signal distributed $N(\theta,\eta_{t-1}^{2})$ (corresponding to the time-(t1)(t-1) distribution of posterior means, conditional on the state) and my belief is about the state is $N(\mu_{it},\eta_{t}^{2})$, then my (state-unconditional) belief about my opponent’s signal is
@@ Appendix VI Payoffs generalizations
[…]
@@
VI.1 Heterogeneous Unknown Initial Shocks
Lemma 10.
Consider the model and payoffs of the main text, with one generalization: $\theta_{i1}:=\theta+\lambda_{i0}(\theta)$, where $\{\lambda_{i0}(\cdot)\}_{i\in\mathcal{I}}$ are bounded, continuous, and increasing in $\theta$ and distributed independently from agents’ information. For all $t\geq 2$, players use identical cutoff strategies.
Proof.
We show this by induction. The inductive step is exactly the same as the inductive step of Lemma 1. For the base case of $t=2$, observe that $a_{i2}=1$ iff
$\mathbb{E}_{i2}[\theta+\lambda_{1}(\theta)]\geq c$
where $i$’s belief is $N(\mu_{i2},\eta_{2}^{2})$. Since $\mu\geq\mu^{\prime}$ implies $N(\mu,\eta_{2}^{2})\geq_{FOSD}N(\mu^{\prime},\eta_{2}^{2})$, it suffices to show that $\lambda_{1}(\theta)$ is increasing. Note that at $t=1$, $a_{i1}=1$ iff
$\mathbb{E}_{i1}[\theta+\lambda_{i0}(\theta)]\geq c\iff\mu_{i1}\geq\mu_{i1}^{*}$
by an analogous argument as above and in Lemma 1, and since $\lambda_{i0}(\theta)$ is increasing. Since $\mu_{i1}$ and $\mu_{i1}^{*}$ are independent random variables on the probability space constructed by Sun and Zhang (2009),
$\lambda_{1}(\theta)=\mathbb{P}\Big{(}\mu_{i1}\geq\mu_{i1}^{*}\Big{)}=\mathbb{E}_{\mu_{i1}^{*}}\Phi\left(\frac{\theta-\mu_{i1}^{*}}{\eta_{t}}\right)$
which is increasing in $\theta$, as desired. ∎
Here, we relax common knowledge about $\lambda_{0}$ in two ways. First, we allow for the initial shock to depend on the underlying fundamental, such that no player knows the value of the initial shock (and therefore must form beliefs about it). Second, we allow for the initial shock to be heterogeneous across players $i$. While this induces heterogeneous cutoffs at $t=1$, this Lemma shows that for all $t\geq 2$, players revert to using identical cutoff strategies.